OpenStudy (anonymous):

The sum of three numbers is 26. The second number is twice the first and the third number is 6 more than the second. Find the numbers. Use the V.E.S.T. system. Can someone explain how to use VEST?

OpenStudy (bibby):

Google says "variable, equation, solve it, test it" Start by writing out the variables. We know that the first variable is some unknown number. What is that?

OpenStudy (anonymous):

The variables would be the first number (x), the second number (2x) and the third number (x+6) ... Right? @bibby

OpenStudy (bibby):

Exactly. Now fir that into the equation "The sum of three numbers is 26. "

OpenStudy (anonymous):

So... x+2x+2x+6=26? @bibby

OpenStudy (bibby):

yes combine like terms

OpenStudy (anonymous):

5x+6=26?

OpenStudy (bibby):

yes

OpenStudy (anonymous):

Okay now I know to take the 6 and subtract. It'd be 5x=20. Then divide: so x=4. But what do I do to get the other two numbers.

OpenStudy (bibby):

you have 3 numbers x 2x 2x+6 and x=4

OpenStudy (anonymous):

x= 4 2x= 8 2x+6= 14 ??? Is that right?

OpenStudy (bibby):

yiss

OpenStudy (anonymous):

Thank you. I thought it was more complex than that. I appreciate it!

OpenStudy (bibby):

no problem